OpenStudy (anonymous):

Determine if the sequence 1, 2, 3, 5, 8, 13... is arithmetic, geometric, or niether.

OpenStudy (anonymous):

this is the fibonacci sequence here's the wikipedia article on it: http://en.wikipedia.org/wiki/Fibonacci_number

OpenStudy (anonymous):

@cnknd so its neither?

OpenStudy (amistre64):

what defines an arith seq? what defines a geo seq?

OpenStudy (anonymous):

I think an arithm seq has constant differences between the terms

OpenStudy (anonymous):

and geo seq has constant ratios

OpenStudy (amistre64):

lets chk for arith then 1-2=-1 2-3=-1 3-5=-2 ..... can be arith then can it

OpenStudy (amistre64):

try geo 2/1 = 2 3/2 = not 2; cant be geo

OpenStudy (anonymous):

arithmetic means (+ or -) so yeah its " arithmetic "

OpenStudy (anonymous):

So it's arith because it's not ratios?

OpenStudy (amistre64):

*cant be arith

OpenStudy (anonymous):

Oh, so it's neither?

OpenStudy (amistre64):

correct, its neither option

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

yw